A 2-horsepower automobile electric starter motor operates at 90 percent efficiency from a 12-V
battery. What is the battery internal resistance if the battery terminal voltage drops to 10 V
when energizing the starter motor?
1
Expert's answer
2021-10-07T09:45:40-0400

2-horsepower:

"2\cdot 746=1492" W

With efficiency 90% motor uses:

"P=1492\/0.9=1658" W

The current:

"I=P\/V=1658\/12=138" A

However that hypothetically loaded motor is not actually running at its rated voltage and power - its at 10 volts. So the current should be scaled by that:

"I=138\cdot 10/12=115" A

So, the battery internal resistance:

"r=\frac{12-10}{115}=0.017\ \Omega"